It is a hot day, and Jacques is thirsty. Here is the value he places on a bottle of water:

4. Problems and Applications Q4

It is a hot day, and Jacques is thirsty. Here is the value he places on a bottle of water:

Value of first bottle:$9
Value of second bottle:$7
Value of third bottle:$4
Value of fourth bottle:$1

From this information, complete the following table by deriving Jacques’s demand schedule.

Price

Quantity Demanded

More than $9
$7.01 to $9
$4.01 to $7
$1.01 to $4
$1 or fewer

Based on Jacques's willingness to pay, plot Jacques's demand curve as a step function on the following graph using blue points (circle symbol) beginning at a quantity of 0 bottles of water.

Jacques's DemandPrice = $6Quantity PurchasedConsumer Surplus012345109876543210Price of WaterQuantity of Water

Suppose the price of a bottle of water is $6.

Use the black line (plus symbol) to draw a price line at $6. Next use the grey point (star symbol) to indicate how many bottles of water Jacques will buy at that price. Finally, use the green point (triangle symbol) to shade the area that represents Jacques's consumer surplus from his purchases.

In this case, Jacques receives in consumer surplus from his water purchase.If the price falls to $3, Jacques now buys bottles of water. This   his consumer surplus to .
Answer #1

In this case Larry receives ……………………consumer surplus from his water purchase.

Consumer surplus = (9-6)*1 + (7-6)*(2-1) = 3+1 =$4

When the price is $3, consumer surplus = (9-3)*1 +(7-3)*(2-1) + (4-3)*(3-2) =6+4+1=$11

Increase in consumer surplus = $11-$4=$7

If the price falls to $3, Larry now buys:- 3 bottles of water

This………….his consumer surplus to………

So correct Answer:- increase; $11